### v1

@@ Abstract
Artificial intelligence (AI) has become a crucial instrument for streamlining processes in various industries, including electrical power systems, as a result of recent digitalization. Algorithms for artificial intelligence are data-driven models that are based on statistical learning theory and are used as a tool to take use of the data that the power system and its users generate. Initially, we perform a thorough literature analysis of artificial intelligence (AI) applications related to renewable energy (RE). Next, we present a thorough analysis of renewable energy factories and assess their suitability, along with a list of the most widely used and appropriate AI algorithms. Nine AI-based strategies are identified here to assist Renewable Energy (RE) in contemporary power systems. This survey paper comprises an extensive review of the several AI techniques used for renewable energy as well as a methodical analysis of the literature for the study of various intelligent system application domains across different disciplines of renewable energy. This literature review identifies the performance and outcomes of nine different research methods by assessing them, and it aims to distill valuable insights into their strengths and limitations. This study also addressed three main topics: using AI technology for renewable power generation, utilizing AI for renewable energy forecasting, and optimizing energy systems. Additionally, it explored AI’s superiority over conventional models in controllability, data handling, cyberattack prevention, smart grid implementation, robotics- AI’s significance in shaping the future of the energy industry. Furthermore, this article outlines future directions in the integration of AI for renewable energy.
@@ 1 Introduction
The fundamental component of financially sound, ecologically responsible, and sustainable electricity generation is renewable energy, or RE. The role that renewable energy plays in reducing climate change [1], [2], [3] is among the most compelling arguments for adopting it. When generating energy, renewable energy sources like solar, wind, and hydroelectric power emit little to no greenhouse gases, in contrast to fossil fuels [4]. When it comes to security and sustainability, renewable energy sources are superior to limited fossil fuels [5]. Natural replenishment of renewable resources, such as sunshine, wind, and water, makes them a steady and dependable energy source free from the price volatility and geopolitical unpredictability of fossil fuels [6]. By supplying power to isolated and neglected areas, renewable energy technologies provide a viable way to combat energy poverty [7]. Off-grid solar systems may provide rural communities with dependable and reasonably priced power, raising living standards and promoting economic growth. Not only is renewable energy crucial, but it is also necessary for a thriving and sustainable future. We can solve a number of issues, including as public health, economic growth, energy security, and climate change, while building a more resilient and sustainable environment for coming generations by embracing renewable energy sources and moving toward a clean energy economy [8], [9].
@@
According to an official report by the International Energy Agency (IEA), the need for fossil fuels to generate power has begun to decline since 2019 [10], coinciding with the rise in the use of renewable energy sources to meet the world’s energy needs. In an effort to improve the efficiency of energy conversion, particularly in terms of RE generation, research on RE technologies is constantly expanding. The current global demand, particularly in developed and developing nations, necessitates the replacement of traditional electrical generation resources like fossil fuels with more sustainable energy technology [11]. Global warming and change in the climate are two harmful environmental problems brought on by fossil fuel-based energy sources. Over the past few decades, there has been an exponential increase in the greenhouse gas emissions from power generation into the environment. In order to tackle the current environmental issue, renewable energy (RE) technologies including solar, wind, hydro, biomass, geothermal, and hydrogen energies have been adopted to create electricity [12]. It is anticipated that in the future, renewable energy sources will provide all of the electricity needed to power residential areas. Data regarding Japan’s 2018 renewable energy percentage is available from the Institute for Sustainable Energy Policies (ISEP) is shown in Figure 1.
@@
International Renewable Energy Agency (IRENA) and Renewable Energy Policy Network for the 21st Century (REN21) [13] provides the important information about Renewable Energy which is recorded in Table 1.
@@
The global pandemic of the corona virus made it difficult to obtain current data on the world’s electrical generation in 2021 from a number of sources, as [10]’s search revealed. Figure 2 illustrates how the difference in energy generation from fossil fuels and renewable energy sources has shrunk between 2010 and 2020. From 121,531 TWh in 2010 to 136,131 TWh in 2019, the amount of fossil fuels used to generate electricity climbed somewhat, but it then declined in 2020. Conversely, the amount of energy generated by renewable sources has grown dramatically, rising from 4098 TWh in 2010 to 7,140 TWh in 2019. In contrast to the production of fossil fuels, renewable energy (RE) accounted for only 3.26–5.60
Figure 2: A comparison between the growth of renewable energy sources and fossil fuel generation (2010–2020) [10]
@@ 1.1 Relevance of Renewable Energy
Renewable energy is extremely important in many aspects that are vital to sustainable development [17]. First and foremost, it plays a critical role in reducing greenhouse gas emissions, since renewable energy sources such as hydroelectric, solar, and wind power emit less greenhouse gases and provide a sustainable substitute for fossil fuels [18].In terms of the economy, the renewable energy sector supports local economies and promotes the shift to a green economy by generating employment, advancing technological innovation, and drawing investment [17]. Furthermore, the utilization of renewable energy improves public health by mitigating air and water pollution that arises from burning fossil fuels, consequently relieving respiratory ailments and other health-related issues. Most importantly, it powers socioeconomic growth by enabling fair access to electricity, especially in isolated and underprivileged areas. In addition, the shift to renewable energy drives technological improvement by promoting energy storage, grid integration, and efficiency research and development [19].
@@ 1.2 Renewable Energy Sector without the influence of Artificial Intelligence
Before the integration of artificial intelligence (AI) into the energy sector, renewable energy technologies such as solar, wind, hydroelectric, geothermal, and biomass were already established and evolving [21]. Developments in these technologies primarily focused on improving efficiency, reducing costs, and addressing environmental concerns. However, despite the advancements brought by AI, there were some limitations and challenges. One of the main lacking aspects was the dependency on traditional methods for data analysis and decision-making, which often lacked real-time insights and predictive capabilities [21]. Additionally, the integration of AI into renewable energy systems required significant investment in infrastructure, data collection, and training algorithms, which posed barriers to adoption, particularly in less developed regions. Moreover, there were concerns regarding data privacy, security, and the potential for AI algorithms to exacerbate biases or inequalities in energy access and distribution. Despite these challenges, the integration of AI has the potential to revolutionize the renewable energy sector by optimizing operations, improving efficiency, and accelerating the transition to a sustainable energy future. At that time researcher used to focus on improvement of renewable energy sources by means of developing tools, energy management system, finding proper solution for those which are harmful for environment. [22] this research article focuses on assessing the present status of electricity access, identifying issues and challenges, and evaluating the effectiveness of renewable energy and energy efficiency in improving electricity access in urban poor areas, specifically in the Korail slum of Dhaka (2017). This research outcome suggests that rooftop solar PV, electricity generation from solid waste, and replacing incandescent lamps with CFLs are promising approaches to enhance electricity access in the Korail slum, with significant potential for surplus electricity generation and cost savings. A separate study [23] examines the impact of environmental regulations on sustainability, taking into account factors such as research and development in renewable energy (RERD), advancements in technology (TI), and the overall economic performance (GDP). The researchers employed wavelet methods, a specialized mathematical tool, to examine the interconnections between these parameters over various time periods and in diverse manners. They discovered that increasing funding for RERD, implementing more stringent environmental regulations, and promoting the adoption of new eco-friendly technology are crucial for safeguarding the environment. Additionally, they discovered significant correlations between CO2 emissions and factors such as technological advancements and economic expansion.
1.3 Artificial Intelligence in the Renewable Energy Sector
Innovations in digital technology have the power to fundamentally alter our energy supply, commerce, and consumption. The artificial intelligence (AI) technology powers the new digitalization paradigm [24]. Intelligent software that maximizes operations and decision-making will autonomously manage the integration of energy supply, demand, and renewable sources into the power grid. AI will be essential to reaching this objective. A lot of data may be gathered by connecting the electricity grid to various devices and sensors. When combined with AI, this data can provide grid operators with fresh perspectives for more effective operation control. It gives energy providers the freedom to deftly modify supply in response to demand. AI can support distributed energy management and microgrid integration [25]. It is challenging to maintain a balance in the energy flow inside the grid when community-scale renewable energy producing units are integrated with the main grid.The quality and congestion problems may be greatly helped by the AI-powered control system. Intelligent Energy Storage (IES) and artificial intelligence combined have the potential to offer the renewable energy sector a dependable and long-lasting solution [25].
The effects of AI on the business and energy sectors are depicted in Figure 5. The influence of AI on energy enterprises across several industries is expected to exceed current expectations [24]. The influence of AI technology on various business kinds over the next five years is covered by the portion of the red line in Figure 5. The vertical axis illustrates the ”Effect of processes,” while the horizontal axis explains the ”Effect of offerings.” The ”Effect of Offerings” offers more chances to experience the influence of AI in many industries (accept or reject as desired), while the ”Effect of Processes” refers to a sequence of events or procedures followed in order to accomplish a certain objective. The majority of firms anticipate greater effects on supply chain management, production and operations, customer-focused activities, and energy information technology (IT).
@@ 1.3 Artificial Intelligence in the Renewable Energy Sector
Data-driven models such as support vector machines (SVMs), k-nearest neighbors (k-NNs), regression trees, multiple linear regression (LR), gradient boosting (GB), and random forest (RF) are commonly used in the field of renewable energy (RE) prediction, leveraging machine learning (ML) techniques. Deep learning models such as long short-term memory (LSTM), gated recurrent units (GRUs), and deep neural networks (DNNs) have been used to predict power consumption and renewable energy (RE) generation across different time periods, yielding outstanding results [27], [28]. In addition, the combination of LSTM with autoencoders (AE) has produced positive outcomes [29].
1.4 Survey Framework
This research focuses on the application of AI methods in the energy industry. In order to enable researchers and readers to evaluate each other’s AI endeavors, goals, fresh, cutting-edge applications, difficulties, and future paths, this study attempts to provide a realistic baseline. Three main topics were covered: (i) utilising current developments in AI technology to create renewable power; (ii) use AI to forecast renewable energy; and (iii) optimizing energy systems. Furthermore, controllability, large data handling, cyberattack prevention, smart grid, IoT, robotics, energy efficiency optimization, predictive maintenance control, and computational efficiency were all areas in which this study examined how AI approaches surpass conventional models. The importance of AI in the next energy industry is another major topic of this study.
@@ 1.4 Survey Framework
The survey structure of this research is organized in a strategic and logical way to cover all the important issues which is shown in the Figure 7. The rest of the paper is organized in this way.
@@ 2.1 Forecasting Renewable Energy Generation using Deep Learning
Authors [30] proposes the ’AB-Net’ architecture, combining autoencoder (AE) and bidirectional long short-term memory (BiLSTM) for one-step forecasting of renewable energy (RE) generation. By acquiring data from various RE sources and deep preprocessing, the model extracts features and forecasts RE generation. Two datasets that are publicly accessible were employed in order to test and assess the efficacy of the suggested method: the wind dataset and the solar dataset. The model consists of three main steps: data acquisition from various renewable energy sources, deep preprocessing to clean and normalize the acquired data, and feature extraction and forecasting using AE and BiLSTM which is shown in Figure 8.
@@
The state computation of a hidden layer in a Long Short-Term Memory (LSTM) network is explained in Equation (1). At each time step $t$, the hidden state $h_{t}$ is updated and maintained based on the previous hidden state $h_{t-1}$ and the current input $x_{t}$ to the layer. In Equation (1), $\sigma_{h}$ denotes the activation function, $W_{xh}$ is the weight matrix from the input to the hidden layer, $W_{hh}$ is the weight matrix for the recurrent connections of the hidden states, and $b_{h}$ is the bias vector for the hidden layer.
Figure 8: The design presented by the authors adheres to a well-defined framework. Data on electricity generation is acquired during Step 1. The second step entails preparing this data. During Step 3, the features are retrieved and then transmitted through the BiLSTM for the purpose of decoding. In the final step, Step 4, the model that has been trained is used to generate predictions. These predictions are then assessed using simple error measures and shown visually through graphs. [30]
AB-Net demonstrated superior performance compared to other models such as BiLSTM, CNN-BiLSTM, and an encoder-decoder (ED) in terms of mean absolute error (MAE), mean squared error (MSE), and root mean square error (RMSE). The solar dataset had mean squared error (MSE), root mean squared error (RMSE), and mean absolute error (MAE) values of 0.0106, 0.1028, and 0.0743, respectively. In contrast, the wind dataset had corresponding values of 0.0004, 0.0189, and 0.0109, respectively. This hybrid network enhances energy management efficiency by facilitating enhanced integration, trading, and control of renewable energy installations.

### v2

@@ Abstract
Artificial intelligence (AI) is becoming an important tool for improving processes in many industries, including electrical power systems, thanks to advancements in digital technology. This paper starts by reviewing how AI is currently used in renewable energy (RE) applications. It then provides an in-depth look at renewable energy systems, assessing their effectiveness and identifying the most commonly used AI algorithms. The review highlights nine AI-based methods that help integrate renewable energy into modern power systems. It also gives a clear overview of AI techniques used in different areas of renewable energy. By analyzing several studies, the paper offers useful insights into the strengths and weaknesses of these methods. Additionally, it focuses on three main areas: how AI is used to generate renewable energy, forecast energy production, and optimize energy systems. The paper also shows how AI outperforms traditional methods in areas like robotics, preventing cyberattacks, processing data, and managing smart grids. The findings show that AI has great potential to improve the energy sector, making it more efficient, secure, and adaptable. The paper concludes by discussing future directions for using AI in renewable energy, highlighting its role in shaping the future of sustainable energy systems.
@@ 1 Introduction
The fundamental component of financially sound, ecologically responsible, and sustainable electricity generation is renewable energy, or RE. The role that renewable energy plays in reducing climate change [1], [2], [3] is among the most compelling arguments for adopting it. Compared to fossil fuels, renewable energy sources such as hydroelectric, solar, and wind power produce minimal to no greenhouse gas emissions throughout the energy generation process [4]. Renewable energy sources outperform finite fossil fuels in terms of sustainability and security [5]. Natural replenishment of renewable resources, such as sunshine, wind, and water, makes them a steady and dependable energy source free from the price volatility and geopolitical unpredictability of fossil fuels [6]. By supplying power to isolated and neglected areas, renewable energy technologies provide a viable way to combat energy poverty [7]. Off-grid solar systems may provide rural communities with dependable and reasonably priced power, raising living standards and promoting economic growth. Not only is renewable energy crucial, but it is also necessary for a thriving and sustainable future. We can solve a number of issues, including as public health, economic growth, energy security, and climate change, while building a more resilient and sustainable environment for coming generations by embracing renewable energy sources and moving toward a clean energy economy [8], [9].
@@
In accordance with the growing adoption of renewable energy sources to fulfill global energy demands, the International Energy Agency’s (IEA) official report states that from 2019 [10], there has been a decrease in the requirement for fossil fuels to come up with power. RE technology research is always growing in an attempt to increase energy conversion efficiency, especially in terms of generation. The contemporary worldwide demand, especially in developed and emerging countries, demands the substitution of environmentally friendly electricity technology for conventional electrical generating resources including fossil fuels [11]. Global warming and change in the climate are two harmful environmental problems brought on by fossil fuel-based energy sources. The amount of greenhouse gases that electricity generation has released into the atmosphere has increased exponentially during the last several decades. Renewable energy (RE) technologies, such as wind, hydroelectricity, solar energy, biomass, geothermal energy, and hydrogen energies, have been utilized to generate power in order to address the present environmental crisis [12]. It is projected that 100% of the electricity required to generate electricity residential areas will come from renewable energy sources in the future. Data regarding Japan’s 2018 renewable energy percentage is available from the Institute for Sustainable Energy Policies (ISEP) is shown in Figure 1.
@@
The crucial data on renewable energy is provided by the International Renewable Energy Agency (IRENA) as well as the Renewable Energy Policy Network over the 21st Century (REN21) [13], and is included in Table 1.
@@
The global pandemic of the corona virus made it difficult to obtain current data on the world’s electrical generation in 2021 from a number of sources, as [10]’s search revealed. The reduction in the production of energy coming from fossil fuels such as and green energy from 2010 and 2020 is seen in Figure 2. The total quantity of fossil fuels utilized to come up with electricity increased considerably, from 121,531 TWh from 2010 to 136,131 TWh starting 2019, before declining in 2020. On the other hand, the quantity of energy produced by green energy has increased significantly, from 4098 TWh by 2010 to 7,140 TWh by 2019. The generation of power from renewable energy (RE) was only 3.26–5.60% between 2010 and 2020, in comparison to the generation of fossil fuels.
Figure 2: An analysis contrasting the production of fossil fuels and renewable energy sources between 2010 and 2020 [10]
@@
This review paper provides the following contributions in the concerned domain. The contributions can be summarized below:
  1. 1.
    Comprehensive Review of AI in Renewable Energy (RE): The paper provides a detailed analysis of how AI is currently applied in the renewable energy sector, offering insights into its use for improving energy efficiency, forecasting, and system optimization.
  2. 2.
    Identification of AI-Based Methods: It highlights nine specific AI-based approaches that support the integration of renewable energy into modern power systems, presenting a clear framework for their application.
  3. 3.
    Comparison of AI Techniques and Traditional Methods: The paper compares the effectiveness of AI techniques with traditional models in areas such as robotics, data processing, and cyberattack prevention, showing how AI provides superior performance.
  4. 4.
    Future Directions for AI Integration: It outlines potential future advancements in the use of AI for renewable energy, focusing on the further optimization of smart grids, energy management, and the development of sustainable power systems.
@@ 1.1 Relevance of Renewable Energy
Numerous factors that are essential to sustainable development depend heavily on renewable energy [17]. Primarily, it is essential for lowering greenhouse gas emissions since renewable energy sources, including wind, solar, and hydroelectric power, produce fewer greenhouse gases and offer a sustainable alternative to fossil fuels [18].Economically speaking, the renewable energy industry stimulates employment, advances technical innovation, and attracts investment, all of which help local economies and the transition to a green economy [17]. Furthermore, the utilization of renewable energy improves public health by mitigating air and water pollution that arises from burning fossil fuels, consequently relieving respiratory ailments and other health-related issues. Most importantly, it powers socioeconomic growth by enabling fair access to electricity, especially in isolated and underprivileged areas. In addition, the shift to renewable energy drives technological improvement by promoting energy storage, grid integration, and efficiency research and development [19].
The major issues can be identified as:
  1. 1.
    Greenhouse Gas Emissions: Renewable energy is crucial for reducing greenhouse gas emissions, as sources like wind, solar, and hydroelectric power produce far fewer emissions compared to fossil fuels [17].
  2. 2.
    Economic Growth: The renewable energy sector promotes job creation, technological innovation, and investment, contributing to local economies and supporting the transition to a green economy [18].
  3. 3.
    Public Health: Utilizing renewable energy helps to mitigate air and water pollution caused by burning fossil fuels, which in turn reduces respiratory and other health issues [17].
  4. 4.
    Socioeconomic Development: Renewable energy enables equitable access to electricity, particularly in remote and underserved areas, fostering socio-economic growth [19].
  5. 5.
    Technological Advancement: The transition to renewable energy drives improvements in energy storage, grid integration, and efficiency, spurring further research and development in these areas [19].
@@ 1.2 Renewable Energy Sector without the influence of Artificial Intelligence
Renewable energy technologies, including solar, wind, hydropower, geothermal, and biomass, were well-established and continuously developing prior to the incorporation technology artificial intelligence (AI) within the energy industry [21]. Developments in these technologies primarily focused on improving efficiency, reducing costs, and addressing environmental concerns. However, despite the advancements brought by AI, there were some limitations and challenges. One of the main lacking aspects was the dependency on traditional methods for data analysis and decision-making, which often lacked real-time insights and predictive capabilities [21]. Adoption was hampered by the additional costs associated with infrastructure, data gathering, and algorithm training that came with integrating AI within renewable energy systems, especially in less developed areas. Moreover, there were concerns regarding data privacy, security, and the potential for AI algorithms to exacerbate biases or inequalities in energy access and distribution. Notwithstanding these difficulties, integrating AI has the power to completely transform the renewable energy industry by streamlining processes, boosting productivity, and hastening the shift to sustainable energy sources. Researchers used to concentrate on improving renewable energy sources at that time by creating tools, energy management systems, and appropriate solutions for those that were environmentally detrimental. In this research article [22], the current state of electricity access is evaluated, problems and obstacles are noted, and the contribution of energy efficiency and renewable energy to the improvement of electricity accessibility in urban poor areas—specifically, the Korail slum of Dhaka—is assessed (2017). This research outcome suggests that rooftop solar PV, electricity generation from solid waste, and replacing incandescent lamps with CFLs are promising approaches to enhance electricity access in the Korail slum, with significant potential for surplus electricity generation and cost savings. A separate study [23] examines the impact of environmental regulations on sustainability, taking into account factors such as research and development in renewable energy (RERD), advancements in technology sector (TI), and the overall economic performance (GDP). The researchers employed wavelet methods, a specialized mathematical tool, to examine the interconnections between these parameters over various time periods and in diverse manners. They discovered that increasing funding for RERD, implementing more stringent environmental regulations, and promoting the adoption of new eco-friendly technology are crucial for safeguarding the environment. Additionally, they discovered significant correlations between CO2 emissions and factors such as technological advancements and economic expansion.
1.3 Artificial Intelligence in the Renewable Power Sector
Digital technology advancements have the potential to drastically change how we use, trade, and supply energy. The new digitalization approach is powered by artificial intelligence (also known as AI) technology [24]. The integration of power demand, supply, and sources of renewable energy into the power grid will be autonomously managed by intelligent software that optimizes operations and decision-making. AI will be necessary to accomplish this goal. By linking different gadgets and sensors to the electrical grid, a vast amount of data could be collected. This data, when paired with AI, can give grid operators new insights for more efficient operation control. It allows energy suppliers the flexibility to skillfully adjust supply in reaction in demand. AI can help with microgrid integration and distributed energy management [25]. Integrating community-scale energy from renewable sources producing units with the main electrical system presents a challenge to maintaining grid balance in energy flow.The AI-powered system of control might potentially significantly alleviate the quality and traffic issues. A trustworthy and long-lasting solution for the renewable energy industry may be provided by the combination of artificial intelligence and intelligent energy storage (IES systems) [25].
Figure 5 shows how AI is affecting the commercial and energy sectors. It is anticipated that AI will have a greater impact on energy firms than currently anticipated in a number of industries [24]. The segment of the red line as shown in Figure 5 represents the impact that AI technology on different business types throughout the next five years. The horizontal axis describes the ”Effect of offerings,” while the vertical axis depicts the ”Effect of processes.” Although the ”Effect of Processes” describes a series of actions taken to achieve a certain goal, the ”Effect of Offerings” gives additional opportunities to experience the effect of AI in numerous sectors (agree to or reject as desired). Most businesses expect increased impacts on customer-focused initiatives, manufacturing and operations, supply chain management, including energy information technology (also known as IT).
@@ 1.3 Artificial Intelligence in the Renewable Power Sector
Data-driven models such as support vector machines (known as SVM), random forest (known as RF), k-nearest neighbors (known as SVM k-NNs), regression trees, multiple linear regression, and gradient boosting (known as GB) are commonly used in the field of renewable energy (RE) prediction, leveraging machine learning (ML) techniques. Deep learning models such as long short-term memory (also known as LSTM), gated recurrent units (also known as GRUs), and deep neural networks (also known as DNNs) have been used to predict consumption of power and renewable energy generation across different time periods, yielding outstanding results [27], [28]. In addition, the combination of LSTM with autoencoders (AE) has produced positive outcomes [29].
1.4 Survey Framework of this Research
The employment of AI techniques in the energy sector is the main topic of this study. This study aims to establish a realistic baseline so that researchers and readers may assess each other’s AI initiatives, ambitions, novel, cutting-edge applications, challenges, and future trajectories. Three main topics were covered: (i) utilising current developments in AI technology to create renewable power; (ii) use AI to forecast renewable energy; and (iii) optimizing energy systems. This research also looked at how AI techniques outperform traditional models in the areas of controllability, huge data handling, cyberattack protection, smart grid, Internet of Things, robotics, conservation of energy optimization, automated maintenance control, as well as computing efficiency. Another important area of this study is the role AI will play in the emerging energy sector.
@@ 1.4 Survey Framework of this Research
The survey structure of this research is organized in a strategic and logical way to cover all the important issues which is shown in the Figure 7. This is how the remaining portion of the paper is structured.
@@ 2 AI Driven Methods For Renewable Energy
AI-driven methods for renewable energy are transforming the way power systems operate by enhancing efficiency and reliability. Machine learning algorithms, such as neural networks and reinforcement learning, are widely used to forecast energy generation, optimize grid management, and improve demand-response systems. These techniques help in accurately predicting renewable energy output based on weather conditions, ensuring better integration of solar and wind power into the grid. AI also plays a key role in energy storage management, optimizing the use of batteries and other energy storage systems. Additionally, AI-driven models contribute to reducing energy waste and improving the overall sustainability of renewable energy systems.
@@ 2.1 Forecasting Renewable Energy Generation using Deep Learning
The authors [30] suggest the ’AB-Net’ architecture allowing one-step modeling of renewable energy (RE) generation, which combines autoencoder (AE) with bidirectional long-short-term (BiLSTM). By acquiring data from various RE sources and deep preprocessing, the model extracts features and forecasts RE generation.The wind database and the solar dataset, two publicly available datasets, were used to test and evaluate the effectiveness of the proposed strategy. The model consists of three main steps: data acquisition from various renewable energy sources, deep preprocessing to clean and normalize the acquired data, and feature extraction and forecasting using AE and BiLSTM which is shown in Figure 8.
@@
Equation (1) describes the state calculation of a hidden layer within LSTM network. The hidden state $h_{t}$ is kept up-to-date and at every interval $t$ depending on the input $x_{t}$ that is currently being fed into the layer as well as the prior hidden state $h_{t-1}$. The activation function is represented by $\sigma_{h}$ in Equation (1), the weight matrix connecting the input to the hidden layer is represented by $W_{xh}$, the weight matrix with regard to recurrent connections of hidden states is represented by $W_{hh}$, and the bias vector for the hidden layer is represented by $b_{h}$.
Figure 8: Design presented by authors adheres to a well-defined framework. Data on electricity generation is acquired during Step 1. The second step entails preparing this data. During third step, features are retrieved as well as then transmitted through BiLSTM in support of purpose of decoding. At the final step, Step 4, model that has been trained is used to generate predictions. These predictions are then assessed using simple error measures and shown visually through graphs. [30]
When compared to other models such an encoder-decoder (ED), CNN-BiLSTM, and BiLSTM, AB-Net performed better with regard of mean absolute error (in short MAE), mean squared error (in short MSE), as well as root mean square error (in short RMSE). The solar dataset’s mean squared error (MSE), mean absolute error (MAE), and root mean square error (RMSE) were 0.0106, 0.1028, along with 0.0743, in that order. The comparable values for the wind dataset were 0.0004, 0.0189, along with 0.0109, in contrast. This hybrid network enhances energy management efficiency by facilitating enhanced integration, trading, and control of renewable energy installations.